OpenStudy (leozap1):

Fan & Medal. A certain business has implemented a new incentive plan whereby an employee earns tokens for every sale they make. For 1 sale, an employee earns 3 token. For 2 sales, an employee earns 4 tokens. For 3 sales, an employee earns 6 tokens. For 4 sales, an employee earns 9 tokens. How many tokens will be earned for eight sales? A. 24 B. 27 C. 31 D. 39

OpenStudy (raffle_snaffle):

look at the token count as we go up from 1 sale to 4 sales

OpenStudy (raffle_snaffle):

do you see a patter?

OpenStudy (raffle_snaffle):

pattern*

OpenStudy (raffle_snaffle):

Do you understand? I am not going to give you the answer. Think about it.

OpenStudy (leozap1):

It goes up 1, then 2, then 3.

OpenStudy (leozap1):

So 9 + 4 = 13 would be #5?

OpenStudy (leozap1):

13 + 5 = 18 for #6

OpenStudy (leozap1):

18 + 6 = 24 for #7

OpenStudy (raffle_snaffle):

yep keep going.

OpenStudy (leozap1):

24 + 7 = 31 for #8

OpenStudy (raffle_snaffle):

there is your answer.